OpenStudy (pokemon23):

When the square of a certain number is diminished by 9times the number the result is 36.Find the number.

OpenStudy (pokemon23):

x\[x^2-9x^2=36\]

OpenStudy (anonymous):

x^2 -9x=36

OpenStudy (anonymous):

is x=12

OpenStudy (dumbcow):

it would be x^2 -9x = 36 x^2 -9x -36=0 (x-12)(x+3) = 0

OpenStudy (pokemon23):

Thanks ;D